Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Help!!!

Expert replies
by manihar.sidharth » Wed Aug 01, 2012 2:58 pm
Please help me with this question And please discuss all the probable cases .Please don't give the the solution using 1-P(Not happening the event)

A fair die has sides labeled with 1, 2, 3, 4, 5, and 6 dots. If the die is rolled 4 times, what is the probability that on at least one roll, the die will show a 6?
A)1/6
b)625/1296
c)649/1296
d)671/1296
e)2/3
OA After some discusion
Join the discussion
Source: — Problem Solving |

by tisrar02 » Wed Aug 01, 2012 5:30 pm
Correct me if i'm wrong but this is what I came up with:

In order to do this question, you NEED to figure out the probability of rolling the die with out getting a 6 first.

Since each case is independent-> (5/6)*(5/6)*(5/6)*(5/6)= 625/1296 which is the probability of NOT rolling a 6

Probability of at least one six is--> 1-(625/1296)= 1296/1296-625/1296

= 671/1296
Join the discussion

by Brent@GMATPrepNow » Thu Aug 02, 2012 6:28 am
manihar.sidharth wrote:Please help me with this question And please discuss all the probable cases .Please don't give the the solution using 1-P(Not happening the event)

A fair die has sides labeled with 1, 2, 3, 4, 5, and 6 dots. If the die is rolled 4 times, what is the probability that on at least one roll, the die will show a 6?
A)1/6
b)625/1296
c)649/1296
d)671/1296
e)2/3
OA After some discusion
The question is testing your ability to solve the question using the complement: P(at least one 6) = 1 - P(getting no 6's)
tisrar02's solution is great, so I won't repeat it here.

To solve the question without using the complement, we need to see that P(at least one 6) = P(exactly one 6) + P(exactly two 6's) + P(exactly three 6's) + P(exactly four 6's).

Lot's of work here!!!

P(exactly one 6)
Consider one case where we get exactly one 6: six on first roll, non-six on second roll, non-six on third roll, and non-six on forth roll.
To make this solution less cumbersome, we'll denote this outcome as {6, ~6, ~6, ~6)
So, P(6, ~6, ~6, ~6) = (1/6)(5/6)(5/6)(5/6)

We also need to consider the outcome {~6, 6, ~6, ~6)
So, P(~6, 6, ~6, ~6) = (5/6)(1/6)(5/6)(5/6)

Notice that we can place the 6 in 4 different places to get exactly one 6, so...
P(exactly one 6) = (4)(1/6)(5/6)(5/6)(5/6) = 500/1296

P(exactly two 6's)
Let's examine one possible outcome: {6, 6, ~6, ~6}
P(6, 6, ~6, ~6) = (1/6)(1/6)(5/6)(5/6)
Of course, {6, 6, ~6, ~6} is just one possible way to get exactly two 6's.
In how many different ways can we get exactly two 6's? Well, in how many different ways can we select 2 places for the 6's?
There are 4 different rolls, and we need to select 2 of them to be 6's. This can be accomplished in 4C2 ways (6 ways).
So, P(exactly two 6's) = (6)(1/6)(1/6)(5/6)(5/6) = 150/1296

P(exactly three 6's)
Let's examine one possible outcome: {6, 6, 6, ~6}
P(6, 6, 6, ~6) = (1/6)(1/6)(1/6)(5/6)
In how many different ways can we select 3 places for the 6's?
There are 4 different rolls, and we need to select 3 of them to be 6's. This can be accomplished in 4C3 ways (4 ways).
So, P(exactly three 6's) = (4)(1/6)(1/6)(1/6)(5/6) = 20/1296

P(exactly four 6's)
There's only one way to accomplish this.
P(6, 6, 6, 6) = (1/6)(1/6)(1/6)(1/6) = 1/1296

So... P(at least one 6) = (500/1296) + (150/1296) + (20/1296) + (1/1296)
= 671/1296
= D

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Brent@GMATPrepNow » Thu Aug 02, 2012 6:31 am
For more information on calculating combinations (like 4C2) in your head, you can watch the following free video: https://www.gmatprepnow.com/module/gmat-counting?id=789

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by manihar.sidharth » Thu Aug 02, 2012 6:40 am
Thanks for such a wonderful explanation Bret and this certainly will help clear the fundamentals but i have a question ..how do we know that we have to consider different outcomes.
I mean to say if I get a six at first position or second position or third position ..how does that matter.
Or Shall i put it this way how shall I come to know that I should be taking the arrangement in the account as well.
Thanks
Sid
Brent@GMATPrepNow wrote:
manihar.sidharth wrote:
P(exactly one 6)
Consider one case where we get exactly one 6: six on first roll, non-six on second roll, non-six on third roll, and non-six on forth roll.
To make this solution less cumbersome, we'll denote this outcome as {6, ~6, ~6, ~6)
So, P(6, ~6, ~6, ~6) = (1/6)(5/6)(5/6)(5/6)

We also need to consider the outcome {~6, 6, ~6, ~6)
So, P(~6, 6, ~6, ~6) = (5/6)(1/6)(5/6)(5/6)

Notice that we can place the 6 in 4 different places to get exactly one 6, so...
P(exactly one 6) = (4)(1/6)(5/6)(5/6)(5/6) = 500/1296

P(exactly two 6's)
Let's examine one possible outcome: {6, 6, ~6, ~6}
P(6, 6, ~6, ~6) = (1/6)(1/6)(5/6)(5/6)
Of course, {6, 6, ~6, ~6} is just one possible way to get exactly two 6's.
In how many different ways can we get exactly two 6's? Well, in how many different ways can we select 2 places for the 6's?
There are 4 different rolls, and we need to select 2 of them to be 6's. This can be accomplished in 4C2 ways (6 ways).
So, P(exactly two 6's) = (6)(1/6)(1/6)(5/6)(5/6) = 150/1296

P(exactly three 6's)
Let's examine one possible outcome: {6, 6, 6, ~6}
P(6, 6, 6, ~6) = (1/6)(1/6)(1/6)(5/6)
In how many different ways can we select 3 places for the 6's?
There are 4 different rolls, and we need to select 3 of them to be 6's. This can be accomplished in 4C3 ways (4 ways).
So, P(exactly three 6's) = (4)(1/6)(1/6)(1/6)(5/6) = 20/1296

P(exactly four 6's)
There's only one way to accomplish this.
P(6, 6, 6, 6) = (1/6)(1/6)(1/6)(1/6) = 1/1296

So... P(at least one 6) = (500/1296) + (150/1296) + (20/1296) + (1/1296)
= 671/1296
= D

Cheers,
Brent
Join the discussion

by Brent@GMATPrepNow » Thu Aug 02, 2012 6:47 am
manihar.sidharth wrote:Thanks for such a wonderful explanation Bret and this certainly will help clear the fundamentals but i have a question ..how do we know that we have to consider different outcomes.
I mean to say if I get a six at first position or second position or third position ..how does that matter.
Or Shall i put it this way how shall I come to know that I should be taking the arrangement in the account as well.
Thanks
Sid
manihar.sidharth wrote:
Hi Sid,

Whenever I'm dealing with the probability of a certain event, like P(exactly one 6), I ask, "What exactly needs to happen in order for this event to occur?"

Here, I see that, to get exactly one six, we can have:
- A 6 on the first roll and no sixes for other rolls (6, ~6, ~6, ~6)
- A non-6, then a 6, then no more sixes (~6, 6, ~6, ~6)
- etc.

If I stop at (6, ~6, ~6, ~6) and calculate only P(6, ~6, ~6, ~6), I'm saying that this is the only way to get exactly one six, and this is not the case.

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by tisrar02 » Thu Aug 02, 2012 6:49 am
In my opinion, when you are dealing with a probability question that states "at least one" of anything, that means you can have 1 or 2 or 3 or whatever. So therefore you want to figure out what the probability of having NONE of the desired outcomes. Because the greatest probability of anything is 1, you would then subtract 1 from the number you received to get the answer you desire.

1-none= at least one

Trying to figure out EVERY case is a lot of work and not recommended on the GMAT.

Hope this helps

Best of luck Sid
Join the discussion

by manihar.sidharth » Thu Aug 02, 2012 6:53 am
Thanks for helping.
Actually i was trying to understand the concept and now I will never make these kinda questions wrong.
Anyways thanks And All the best for you test!!!
tisrar02 wrote:In my opinion, when you are dealing with a probability question that states "at least one" of anything, that means you can have 1 or 2 or 3 or whatever. So therefore you want to figure out what the probability of having NONE of the desired outcomes. Because the greatest probability of anything is 1, you would then subtract 1 from the number you received to get the answer you desire.

1-none= at least one

Trying to figure out EVERY case is a lot of work and not recommended on the GMAT.

Hope this helps

Best of luck Sid
Join the discussion

by Brent@GMATPrepNow » Thu Aug 02, 2012 6:54 am
tisrar02 wrote:In my opinion, when you are dealing with a probability question that states "at least one" of anything, that means you can have 1 or 2 or 3 or whatever. So therefore you want to figure out what the probability of having NONE of the desired outcomes. Because the greatest probability of anything is 1, you would then subtract 1 from the number you received to get the answer you desire.

1-none= at least one

Trying to figure out EVERY case is a lot of work and not recommended on the GMAT.

Hope this helps

Best of luck Sid
Agreed!
That said, if someone failed to use the complement for this question, they would need to solve it using a different approach.

Alternatively, the question could have asked for something like P(exactly two 6's), in which case, we'd need to use the method described above.

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by tisrar02 » Thu Aug 02, 2012 6:59 am
Brent@GMATPrepNow wrote:
tisrar02 wrote:In my opinion, when you are dealing with a probability question that states "at least one" of anything, that means you can have 1 or 2 or 3 or whatever. So therefore you want to figure out what the probability of having NONE of the desired outcomes. Because the greatest probability of anything is 1, you would then subtract 1 from the number you received to get the answer you desire.

1-none= at least one

Trying to figure out EVERY case is a lot of work and not recommended on the GMAT.

Hope this helps

Best of luck Sid
Agreed!
That said, if someone failed to use the complement for this question, they would need to solve it using a different approach.

Alternatively, the question could have asked for something like P(exactly two 6's), in which case, we'd need to use the method described above.

Cheers,
Brent
Figured it out. Thanks Brent for the alternative view!
Join the discussion